# Deterministic Dynamic Edge-Colouring

### v1

@@ 6.4 Analysing the data structures
Lemma 60.
Let $f$ be marked with $j\leq i$. Then for any $w$, we can update $\mathcal{D}_{i}(f,w)$ in $\tilde{O}(3^{\sqrt{\log n}}(\Delta_{\max}+1)^{3\sqrt{\log n}+4}\cdot{}\ell^{3}\cdot{}\log^{2}{n})$ time.
Proof.
Note firstly that the $\tilde{O}$-notation takes care of any $\log n$ factors arising from inserting, searching or deleting things from BBST’s. Here it is important to note that the BBST’s used to update the data structures never have sizes greater than $\tilde{O}(\ell^{2}\cdot{}\log{n}\cdot{}3^{\sqrt{\log n}}(\Delta_{\max}+1)^{3\sqrt{\log n}+2})$. Indeed, that is the maximum number of chain extensions consistent with Lemma 13, stored in the neighbourhoods of the first $\ell$ vertices along $P_{1}$.
Secondly, Lemma 59 implies that at most $\sqrt{\log n}\cdot{}3^{\sqrt{\log n}}(\Delta_{\max}+1)^{3\sqrt{\log n}+2}\cdot{}\ell$ of these extensions can be overlapping with the first-step of the process in the sense of Definition 15.
Hence, at most $3^{\sqrt{\log n}}(\Delta_{\max}+1)^{3\sqrt{\log n}+2}\cdot{}\ell\cdot{}\log n$ such chains need to be checked, before $\mathcal{D}_{i}(f,w)$ is either full or contains all semi-consistent and augmenting $i$-step processes. Each extension can be checked in $\tilde{O}((i+1)\cdot{}\ell\cdot{}(\Delta_{\max}+1)^{3})$ time by Lemma 58 and since it only takes $\tilde{O}(\ell\Delta_{\max}^{3})$ to compute and locate the correct choices of $\mathcal{D}_{i-1}(\cdot{},\cdot{})$ to look for extension suggestions in. By noting that $i\leq\sqrt{\log n}$, the Lemma follows. ∎
Combining all of the above finalises the proof of Theorem 51.
6.5 The proof of Theorem 5
In this section, we present a proof of Theorem 5 (restated below for convenience).
Theorem 61 (Identical to Theorem 5).
@@ 6.5 The proof of Theorem 5
[…]
@@
  • Then we pick an arbitrary vertex $u\in e$, and construct a consistent fan according to Lemma 12. The fan ends at a vertex $y_{1}$ and outputs a bichromatic path $P$.
  • If $|P|\leq\ell$, we have discovered an augmenting chain, and we can immediately jump to step 5)5)5 ). Otherwise, we look in $\mathcal{D}_{i}(r(y_{1})r(P),u)$ for all $i$ in order to find a consistent and augmenting $j$-step process $C$ for some $j$. Note that Corollary 46 and Theorem 51 guarantees that such a chain exists.
  • According to Lemma 18, $C$ guarantees some augmenting and consistent $j$-step Vizing chain. We can explicitly compute this chain by tracing the process and computing the consistent fans according to Lemma 13.
[…]
  • Finally since the chain is augmenting, after shifting it the remaining un-coloured edge $f$ has a free colour $\kappa$. We colour it using $\texttt{colour}(f,\kappa)$.
@@
[…]
@@
[…]
@@ 7 Putting it all together
In this Section, we combine the results of the previous section with a degree-scheduler from [DHZ19] to show Theorem 2
Theorem 62 (Identical to Theorem 2).
Let $G$ be a dynamic graph, and $10<\varepsilon<$ 1 a given parameter. We can maintain a proper $((1+\varepsilon)\Delta)$-edge-colouring of $G$ in $2^{\tilde{O}_{\log\varepsilon^{-1}}(\sqrt{\log n})}$ amortised update time per operation.
@@ 7.1 The final algorithm
For now, we will still assume that we have access to some upper bound $\Delta_{\max}$ on the maximum degree of $G$ throughout the entire update sequence. We will then get rid of this assumption in the next section by using a degree-scheduler due to Duan, He, and Zhang [DHZ19].
The algorithm is now the following: If $\Delta_{\max}\leq\tfrac{10^{7}\log^{5}n}{\varepsilon^{3}}$: we simply apply Theorem 5 directly. Since $\Delta_{\max}+1\leq\lceil(1+\varepsilon)\Delta_{\max}\rceil$ for any $\varepsilon>0$, we are immediately done by Theorem 5.
Otherwise, if $\Delta_{\max}>\tfrac{10^{7}\log^{5}n}{\varepsilon^{3}}$, we first apply Theorem 4 to get a degree-splitting hierarchy of height $0<h\leq 2\sqrt{\log n}$. Then we apply Theorem 5 with the upper bound on the maximum degree equal to $\Delta^{\prime}_{\max}$, where $\Delta^{\prime}_{\max}=\hat{\Delta}_{h}$, to each graph of $\mathcal{G}_{h}$, using a distinct palette to colour each graph. Correctness of the algorithm then follows immediately from Theorem 4 and Theorem 5.
As for the update time of the algorithm, we note that by Theorem 4 each sequence of $\sigma$ updates to $G$ results in at most $\sigma\cdot{}2^{\tilde{O}_{\log\varepsilon^{-1}}(\sqrt{\log n})}$ updates to graphs in $\mathcal{G}_{h}$. We can handle each such update in $2^{\tilde{O}_{\log\varepsilon^{-1}}(\sqrt{\log n})}$ time by Theorem 5, so in total, we process the $\sigma$ updates in $\sigma\cdot{}2^{\tilde{O}_{\log\varepsilon^{-1}}(\sqrt{\log n})}$ time, thus resulting in an amortised update time of $2^{\tilde{O}_{\log\varepsilon^{-1}}(\sqrt{\log n})}$ time, as claimed.
Finally, we analyse the number of colours used. We note that by Theorem 4, there exists non-negative integers $i,j$ with $i+j=h$ such that:
@@
$|\mathcal{G}_{h}|\cdot{}((1+\tfrac{\varepsilon}{16})\frac{\Delta_{\max}}{t_{1}^{i}t_{2}^{j}})+1)\leq(1+\tfrac{\varepsilon}{2})\Delta_{max}+t_{1}^{i}t_{2}^{j}\leq(1+\varepsilon)\Delta_{\max}$
Here we used the fact that Theorem 4 terminates with a hierarchy for which $\Delta_{\max}\geq\tfrac{2\cdot{}t_{1}^{i}t_{2}^{j}}{\varepsilon}$.
@@ 7.2 Dynamic $\Delta$
In this subsection, we finish the proof of Theorem 2. This last part contains no new ideas, as we can use the degree-scheduler of Duan, He and Chang [DHZ19] essentially ad verbatim. For the sake of completeness, we briefly sketch their approach below.
Let $\mu$ be some constant not specified until later (we will ultimately set $\mu=\Theta(\varepsilon)$). Consider the exponentially increasing intervals $I_{i}:=[(1+\mu)^{i-1},(1+\mu)^{i}]$ for $i\in\{0,1,\dots,s\}$ with $s$ minimum subject to $(1+\mu)^{s}\geq n$. For each $i$, we will maintain a subgraph of $G$, say $G_{i}$ with the guarantees that if $\Delta(G)\in I_{i}$, then $G_{i}=G$, and that for all $i$, we have $\Delta(G_{i})\leq(1+\mu)^{i+1}$ always. Furthermore, we will at all times maintain some index $j$ for which $\Delta(G)\in I_{j}$.
Given the above, one can run any $(1+\varepsilon^{\prime})\Delta_{\max}$-edge-colouring algorithm on all of the $G_{i}$’s, and then let the colouring of $G_{j}$ be the currently valid edge-colouring. If one sets $\Delta_{\max,i}=(1+\mu)^{i+1}$ in the instance of the edge-colouring algorithm on $G_{i}$, one has the guarantee that at all times the algorithm maintains a $(1+\mu)^{2}(1+\varepsilon^{\prime})\Delta$-edge-colouring. Indeed, at all times in the copy currently indexed by $j$ will have $\Delta\cdot{}(1+\mu)^{2}\geq\Delta_{\max,j}$. Setting $\varepsilon^{\prime},\mu\leq\tfrac{\varepsilon}{3}$ is then sufficient to achieve the correct number of colours.
In order to maintain this information, Duan, He and Chang do as follows: let $\delta(e)$ be the maximum degree of an endpoint of $e$. If $\delta(e)\leq(1+\mu)^{i}$, we add $e$ to $G_{i}$. Now $e$ stays in $G_{i}$ until $\delta(e)>(1+\mu)^{i+1}$, at which point in time it is removed, and not added back until $\delta(e)\leq(1+\mu)^{i}$ again. It is now easy to check that that every time some vertex $w$ crosses the threshold $(1+\mu)^{k}$ for some $k$, it performs at most $deg(w)$ insertions into $G_{k}$ or at most $deg(w)$ deletions from $G_{k}$. Between any set of insertions into some $G_{k}$, $deg(w)$ must have increased beyond $(1+\mu)^{k+1}$ and back down to $(1+\mu)^{k}$. All of the deletions incident to $w$ that decreased the degree from $(1+\mu)^{k+1}$ to $(1+\mu)^{k}$ can be charged the cost of performing the $deg(w)$ insertions into $G_{k}$. Deletions can be analysed in a symmetric way. In total, the amortised recourse of performing these changes is upper bounded by $O(\mu^{-1})$ for each update. Since there are at most $O(\tfrac{\log n}{\mu})$ choices of $i$, the total amortised recourse and update time becomes $O(\tfrac{\log n}{\mu})$. One can easily keep track of $j$ or directly calculate $j$ using knowledge of the current maximum degree $\Delta$. Therefore picking $\mu,\varepsilon^{\prime}=\theta(\varepsilon)$ with $\mu,\varepsilon^{\prime}\leq\tfrac{\varepsilon}{3}$ yields the corollary.
@@ 8 Acknowledgements
The author would like to thank Eva Rotenberg for the support and encouragements.

### v2

@@ 6.4 Analysing the data structures
Lemma 6.30.
Let $f$ be marked with $j\leq i$. Then for any $w$, we can update $\mathcal{D}_{i}(f,w)$ in $\tilde{O}(3^{\sqrt{\log n}}(\Delta_{\max}+1)^{3\sqrt{\log n}+4}\cdot{}\ell\cdot{}a^{2}\cdot{}\log^{2}{n})$ time.
Proof 38.
Note firstly that the $\tilde{O}-$notation takes care of any $\log n$ factors arising from inserting, searching or deleting things from BBST’s.
We need to check at most $\ell$ extensions $x$. Each can be checked in $\tilde{O}(\sqrt{\log n}\cdot{}\ell\cdot{}a^{2}\cdot{}(\Delta_{\max}+1)^{3})$ time by Lemma 6.29. After checking these extensions, we can create the set of strings to store in $\mathcal{D}_{i}(f,w)$ in $O(\tfrac{a}{2})$ time by storing at least one non-blocked string for each non-blocked extension.
Picking $\ell=10^{4}(10\Delta_{\max}+1)^{50\sqrt{\log n}}$ (any slightly larger value also works) and $\tfrac{\ell}{(2^{3\sqrt{\log n}}(\Delta_{\max}+1)^{8\sqrt{\log n}})}\leq a\leq\tfrac{\ell}{(2^{3\sqrt{\log n}}(\Delta_{\max}+1)^{6\sqrt{\log n}})}$, and combining all of the above finalizes the proof of Theorem 6.23.
6.5 The proof of Theorem 1.5
In this section, we present a proof of Theorem 1.5 (restated below for convenience).
Theorem 6.31 (Identical to Theorem 1.5).
@@
Firstly, fix parameters $\ell=10^{4}(10\Delta_{\max}+1)^{50\sqrt{\log n}}$ and any value of $a$ satisfying that
$\tfrac{\ell}{(2^{3\sqrt{\log n}}(\Delta_{\max}+1)^{8\sqrt{\log n}})}\leq a\leq\tfrac{\ell}{(2^{3\sqrt{\log n}}(\Delta_{\max}+1)^{6\sqrt{\log n}})}$ so that we may apply the lemmas from the earlier parts of the section.
@@
Maintaining the data structures from Section 6.3 represents most of the work required to show Theorem 1.5. The idea is the following. We will indeed maintain these data structures and ensure that after every update terminates, the colouring $c$ maintained by the algorithm of Theorem 6.23 is a proper $(\Delta_{\max}+1)$-edge-colouring. Given a call to $\texttt{delete}(e)$, we first call $\texttt{un-colour}(e)$ and then subsequently $\texttt{remove}(e)$ in the previous data structures and in our representation of $G$. Given a call to $\texttt{insert}(e)$, we do the following steps:
@@
  • Then we pick an arbitrary vertex $u\in e$, and construct a consistent fan according to Lemma 3.4. The fan ends at a vertex $y_{1}$ and outputs a bichromatic path $P$.
  • If $|P|\leq\ell+1$, we have discovered an augmenting chain, and we can immediately jump to step $5)$. Otherwise, we look in $\mathcal{D}_{\sqrt{\log n}}(r(y_{1})r(P),u)$ to find an augmenting and consistent stepping process. By Theorem 6.23 this set is non-empty.
  • Then we convert this string into the stepping process it presents and then into a Vizing chain in its pre-image. By Lemma 3.10 this Vizing chain is augmenting.
  • We shift the colouring $c$ according to this chain using the operations $\texttt{un-colour}(\cdot{})$ and $\texttt{colour}(\cdot{},\cdot{})$. Since the colouring is shifted along a chain, we ensure that $c$ is at all times proper, and so all of the required assumptions hold.
  • Finally, since the chain is augmenting, after shifting it, the remaining un-coloured edge $f$ has a free colour $\kappa$. We colour it using $\texttt{colour}(f,\kappa)$.
@@
Theorem 6.5 guarantees the existence of a consistent and augmenting Vizing chain of the form, we seek. Corollary 6.13 guarantees that the data structure $\mathcal{D}_{\sqrt{\log n}}(r(y_{1})r(P),u)$ contains at least one string of the claimed form. Next, applying Lemma 3.10 ensures that this process guarantees a consistent and augmenting Vizing chain. Finally, correctness follows by Theorem 6.23, as the Theorem guarantees the correctness of $\texttt{un-colour}(\cdot{})$ and $\texttt{colour}(\cdot{},\cdot{})$.
@@
The running time of $\texttt{delete}(e)$ follows immediately from Theorem 6.23. As for $\texttt{insert}(e)$, we note that since the shiftable-chain has length $\tilde{O}(\sqrt{\log n}\ell)$, we need only call $\texttt{un-colour}(\cdot{})$ and $\texttt{colour}(\cdot{},\cdot{})$ $\tilde{O}(\sqrt{\log n}(\ell+(\Delta_{\max}))$ times. This part of the process dominates the update time, as we can trace the process in $\tilde{O}(\sqrt{\log n}\cdot{}\ell)$ time, and compute the at most $\sqrt{\log n}$ fans in $\tilde{O}(\Delta^{3}_{\max})$ time each. This concludes the proof.
@@ 7 Putting it all together
In this Section, we combine the results of the previous section with a degree-scheduler from [duan] to show Theorem 1.2
Theorem 7.1 (Identical to Theorem 1.2).
Let $G$ be a dynamic graph, and $0<\varepsilon<1$ a given parameter. We can maintain a proper $(1+\varepsilon)\Delta$-edge-colouring of $G$ in $2^{\tilde{O}_{\log\varepsilon^{-1}}(\sqrt{\log n})}$ amortised update time per operation.
@@ 7.1 The final algorithm
For now, we will still assume that we have access to some upper bound $\Delta_{\max}$ on the maximum degree of $G$ throughout the entire update sequence. We will then get rid of this assumption in the next subsection by using a degree-scheduler due to Duan, He, and Zhang [duan].
The algorithm is now the following: If $\Delta_{\max}\leq\tfrac{10^{7}\log^{5}n}{\varepsilon^{3}}$: we simply apply Theorem 1.5 directly. Since $\Delta_{\max}+1\leq\lceil(1+\varepsilon)\Delta_{\max}\rceil$ for any $\varepsilon>0$, we are immediately done by Theorem 1.5.
Otherwise, if $\Delta_{\max}>\tfrac{10^{7}\log^{5}n}{\varepsilon^{3}}$, we first apply Theorem 1.4 to get a degree-splitting hierarchy of height $0<h\leq 4\sqrt{\log n}$. Then we apply Theorem 1.5 with the upper bound on the maximum degree equal to $\Delta^{\prime}_{\max}$, where $\Delta^{\prime}_{\max}=\hat{\Delta}_{h}$, to each graph of $\mathcal{G}_{h}$, using a distinct palette to colour each graph. Correctness of the algorithm then follows immediately from Theorem 1.4 and Theorem 1.5.
As for the update time of the algorithm, we note that by Theorem 1.4 each sequence of $\sigma$ updates to $G$ results in at most $\sigma\cdot{}2^{\tilde{O}_{\log\varepsilon^{-1}}(\sqrt{\log n})}$ updates to graphs in $\mathcal{G}_{h}$ and at most $2^{\tilde{O}_{\log\varepsilon^{-1}}(\sqrt{\log n})}$ total update time. We can handle each such update to a graph in $\mathcal{G}_{h}$ in $2^{\tilde{O}_{\log\varepsilon^{-1}}(\sqrt{\log n})}$ time by Theorem 1.5, so in total, we process the $\sigma$ updates in $\sigma\cdot{}2^{\tilde{O}_{\log\varepsilon^{-1}}(\sqrt{\log n})}$ time, thus resulting in an amortised update time of $2^{\tilde{O}_{\log\varepsilon^{-1}}(\sqrt{\log n})}$ time, as claimed.
Finally, we analyse the number of colours used. We note that by Theorem 1.4, there exists non-negative integers $i,j$ with $i+j=h$ such that:
@@
$|\mathcal{G}_{h}|\cdot{}((1+\tfrac{\varepsilon}{16})\frac{\Delta_{\max}}{t_{1}^{i}t_{2}^{j}}+1)\leq(1+\tfrac{\varepsilon}{2})\Delta_{\max}+t_{1}^{i}t_{2}^{j}\leq(1+\varepsilon)\Delta_{\max}$
Here we used the fact that Theorem 1.4 terminates with a hierarchy for which $\Delta_{\max}\geq\tfrac{2\cdot{}t_{1}^{i}t_{2}^{j}}{\varepsilon}$.
@@ 7.2 Dynamic $\Delta$
In this subsection, we finish the proof of Theorem 1.2. This last part contains no new ideas, as we can use the degree-scheduler of Duan, He and Chang [duan] essentially ad verbatim. For the sake of completeness, we briefly sketch their approach below.
Let $\mu$ be some constant not specified until later (we will ultimately set $\mu=\Theta(\varepsilon)$). Consider the exponentially increasing intervals $I_{i}:=[(1+\mu)^{i-1},(1+\mu)^{i}]$ for $i\in\{1,\dots,s\}$ with $s$ minimum subject to $(1+\mu)^{s}\geq n$. For each $i$, we will maintain a subgraph of $G$, say $G_{i}$ with the guarantees that if $\Delta(G)\in I_{i}$, then $G_{i}=G$, and that for all $i$, we have $\Delta(G_{i})\leq(1+\mu)^{i+1}$ always. Furthermore, we will at all times maintain some index $j$ for which $\Delta(G)\in I_{j}$.
Given the above, one can run any $(1+\varepsilon^{\prime})\Delta_{\max}$-edge-colouring algorithm on all of the $G_{i}$’s, and then let the colouring of $G_{j}$ be the currently valid edge-colouring. If one sets $\Delta_{\max,i}=(1+\mu)^{i+1}$ in the instance of the edge-colouring algorithm on $G_{i}$, one has the guarantee that at all times the algorithm maintains a $(1+\mu)^{2}(1+\varepsilon^{\prime})\Delta$-edge-colouring. Indeed, at all times in the copy currently indexed by $j$ will have $\Delta\cdot{}(1+\mu)^{2}\geq\Delta_{\max,j}$. Setting $\varepsilon^{\prime},\mu\leq\tfrac{\varepsilon}{7}$ is then sufficient to achieve the correct number of colours.
In order to maintain this information, Duan, He and Chang do as follows: let $\delta(e)$ be the maximum degree of an endpoint of $e$. If $\delta(e)\leq(1+\mu)^{i}$, we add $e$ to $G_{i}$. Now $e$ stays in $G_{i}$ until $\delta(e)>(1+\mu)^{i+1}$, at which point in time it is removed, and not added back until $\delta(e)\leq(1+\mu)^{i}$ again. It is now easy to check that that every time some vertex $w$ crosses the threshold $(1+\mu)^{k}$ for some $k$, it performs at most $deg(w)$ insertions into $G_{k}$ or at most $deg(w)$ deletions from $G_{k-1}$. Between any set of insertions into some $G_{k}$, $deg(w)$ must have increased beyond $(1+\mu)^{k+1}$ and back down to $(1+\mu)^{k}$. All of the deletions incident to $w$ that decreased the degree from $(1+\mu)^{k+1}$ to $(1+\mu)^{k}$ can be charged the cost of performing the $deg(w)$ insertions into $G_{k}$. Deletions can be analysed in a symmetric way. In total, the amortised recourse of performing these changes is upper bounded by $O(\mu^{-1})$ for each update. Since there are at most $O(\tfrac{\log n}{\mu})$ choices of $i$, the total amortised recourse and update time becomes $O(\tfrac{\log n}{\mu})$. One can easily keep track of $j$ or directly calculate $j$ using knowledge of the current maximum degree $\Delta$. Therefore picking $\mu,\varepsilon^{\prime}=\theta(\varepsilon)$ with $\mu,\varepsilon^{\prime}\leq\tfrac{\varepsilon}{7}$ yields the corollary.
@@ 8 Acknowledgements
The author would like to thank Eva Rotenberg for the support, help, and encouragements.